News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 20 July 2012 9.00am EDT

During the last week the particle radiation due to Solar activity following the Jul 12 safing event subsided enough to resume observing. Loads for a replanned schedule were uplinked to the observatory on Jul 13 with 64.7ks of science loss.

The observing schedule was interrupted again due to high radiation associated with solar activity. The loads were halted at 5:07pm EDT on Jul 14 through a ground command to execute the SI Safing SCS 107. This ensured that the accumulated radiation dose for ACIS remained below the allowed threshold. All spacecraft actions were nominal. Real-time procedures were executed on Jul 15 to disable SCS-29, which was enabled during the safing sequence, and to start a collection of ACIS CCD charge-transfer inefficiency data. A replanned schedule for resuming science observations was uplinked on Jul 15 with 80.2ks of science loss.

The observing schedule was again interrupted by a high-radiation event that activated the science instrument safing sequence (SCS 107) and stopped observing at 7:44am EDT on Jul 19. All spacecraft actions were nominal. A real-time procedure was executed on Jul 19 to disable SCS-29, which was enabled during the safing sequence. Loads for a replanned schedule were uplinked to the observatory on Jul 19 with 56.5ks of science loss.

A real-time procedure was executed on Jul 19 to update the on-board ephemeris.

The schedule of targets for the replan and the next week is shown below and includes an observation of GRB 120711A, which was accepted as a Target of Opportunity on Jul 17, an observation of Holmberg IX X-1 coordinated with HST, and observations of Sgr A* coordinated with NuSTAR, Keck, VLT, SMA, CSO, CARMA, ALMA, EHT, HESS, and ATCA.



------------------------------------------
HolmbergIXX-1          ACIS-S       Jul 20
       M31                    ACIS-I
       J1458.4-2447           ACIS-I
       SgrA*                  ACIS-S/HETG  Jul 20
       Radiation Belts
       SgrA*                  ACIS-S/HETG
       Radiation Belts                     Jul 23
       GRB120711A             ACIS-S       Jul 24
       SgrA*                  ACIS-S/HETG
       Radiation Belts                     Jul 26
       SgrA*                  ACIS-S/HETG
       PSRJ1548-4821          ACIS-S       Jul 28
       B2224+65               ACIS-S
       Radiation Belts                     Jul 29
       MAXIJ0556-332          ACIS-S
       B2224+65               ACIS-S
       SWIFTJ1749.4-2807      ACIS-S       Jul 30
       Terzan5                ACIS-S

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports